one cup of unpopped popcorn will result in two quarts of popped popcorn. how many cups of unpopped popcorn are needed to produce 24 quarts of popped popcorn?
Answer by checkley77(12844)

1/2=x/24 cross multiply.
2x=24
x=24/2
x=12 cups od unpopped popcorn is needed.